# ARFlow: Human Action-Reaction Flow Matching with Physical Guidance

### v2

@@ Abstract
Human action-reaction synthesis, a fundamental challenge in modeling causal human interactions, plays a critical role in applications ranging from virtual reality to social robotics. While diffusion-based models have demonstrated promising performance, they exhibit two key limitations for interaction synthesis: reliance on complex noise-to-reaction generators with intricate conditional mechanisms, and frequent physical violations in generated motions. To address these issues, we propose Action-Reaction Flow Matching (ARFlow), a novel framework that establishes direct action-to-reaction mappings, eliminating the need for complex conditional mechanisms. Our approach introduces two key innovations: an $\mathbf{x}_{1}$-prediction method that directly outputs human motions instead of velocity fields, enabling explicit constraint enforcement; and a training-free, gradient-based physical guidance mechanism that effectively prevents body penetration artifacts during sampling. Extensive experiments on NTU120 and Chi3D datasets demonstrate that ARFlow not only outperforms existing methods in terms of Fréchet Inception Distance and motion diversity but also significantly reduces body collisions, as measured by our new Intersection Volume and Intersection Frequency metrics.
$\dagger$\$$\dagger$\$footnotetext: Corresponding authors.
@@ 1 Introduction
To address these challenges, we propose Action-Reaction Flow Matching (ARFlow), a novel framework that fundamentally resolves these limitations. Unlike diffusion models constrained by noise-data mappings, flow matching [38, 42, 1, 45] naturally models paired distributions through linear interpolation between endpoints (See Fig. 1), making it inherently suitable for action-reaction synthesis. ARFlow establishes direct pathways between action and reaction distributions, enabling simpler training and faster inference compared to diffusion-based methods. Instead of traditional vector field prediction ($\mathbf{v}$-prediction) in standard flow matching [38, 25], ARFlow directly outputs clean body poses ($\mathbf{x}_{1}$-prediction). This innovation allows direct application of geometric constraints during training and better alignment with motion generation objectives. We further develop a specialized sampling algorithm optimized for $\mathbf{x}_{1}$-prediction in human motion generation. To eliminate unrealistic body collisions between characters, we introduce a penetration-aware sampling process that maintains physical plausibility through gradient guidance without compromising motion quality. Additionally, we introduce two new physics-aware evaluation metrics: Intersection Volume (IV) and Intersection Frequency (IF). Our main contributions are threefold:
@@
  • We propose a new $\mathbf{x}_{1}$-prediction method under flow matching specifically designed to integrate physical constraints during both training and sampling. This combined approach successfully prevents body collisions between characters while preserving natural movement.
@@
Figure 2: Pipeline of ARFlow. (a) For a sampled timestep $t$, we linearly interpolate a coupled action-reaction pair as Eq. 4 to produce the intermediate state $x^{1:H}_{t}$, which is then turns into a d-dimensional latent feature through a linear layer. We use Transformer Decoder Units to directly predict clean reaction motions. (b) After training the networks in (a), our ARFlow uses them for $\mathbf{x}_{1}$-$prediction$ based sampling. The sampling process is further guided by the gradient of $\mathcal{L}_{pene}$ to generate physically plausible reactions.
@@ 2.2 Human Action-Reaction Synthesis
Different from human-human interaction [67, 37, 55, 26, 57, 64], human action-reaction synthesis is causal and asymmetric [69, 40, 68]. To address this task, researchers have leveraged large language models [54, 56, 28], VAE-based methods [10, 43, 44]. However, these methods cannot capture fine-grained representations and ensure diversity, and diffusion-based methods [36, 57] are limited to the “offline” and “constrained” setting of human reaction generation, failing to generate instant and intention agnostic reactions. More recently, ReGenNet [69] introduce a diffusion-based transformer decoder framework and treat action sequence as conditional signal for online reaction generation. However, it often produce physically-implausible inter-penetrations between the actor and reactor since they disregard physical constraints in the generative process. Our method addresses this problem by guiding the sampling process with physical constraint guidance.
@@ 2.3 Flow Matching
Flow Matching [38, 42, 1, 45] has emerged as an efficient alternative to diffusion models, offering linear generation trajectories through ODE solvers. This paradigm enables simplified training and accelerated inference [16, 39], with successful applications spanning images [38, 16], audio [34], video [5], and point clouds [65]. In motion generation, MotionFlow [25] demonstrates comparable performance to diffusion models with faster sampling. Notably, Flow Matching inherently models transitions between arbitrary distributions through transport maps, making it particularly suitable for paired data modeling. Despite these advantages, its potential for action-reaction synthesis remains unexplored. Our work bridges this gap by establishing direct action-to-reaction mappings without complex conditional mechanisms.
@@ 3.1 Human Action-Reaction Flow Matching
${\mathcal{L}_{\text{fm}}=\mathbb{E}_{\mathbf{x}_{1}\sim q(\mathbf{x}_{1}),\mathbf{x}_{0}\sim p(\mathbf{x}_{0}),t\sim[0,1]}[\|\mathbf{x}_{1}-G_{\theta}(\mathbf{x}_{t},t,c)\|_{2}^{2}].}$ (5)
@@
Since our neural network outputs $\hat{\mathbf{x}}_{1}$, we require to construct an equivalent relationship between the neural network’s predictions of $\mathbf{v}$ and $\mathbf{x}_{1}$. The equivalent form of parameterization Equation 21 derived from our appendix is as follows:
Then, our sampling based on $\mathbf{x}_{1}$-prediction can be achieved by first sampling $\mathbf{x}_{0}$ and then solving Equation 1 employing an ODE solver [52, 33, 2] through our trained neural network $\hat{\mathbf{x}}_{1}=G_{\theta}(\mathbf{x}_{t},t,c)$. We use the Euler ODE solver and discretization process involves dividing the procedure into $N$ steps, leading to the following formulation:
@@
where the integer time step $t_{1}=0<t_{2}<\cdots<t_{N}=1$. By using equivalent form of parameterization Equation 7, we finally obtain our flow maching sampling formulation based on $\mathbf{x}_{1}$-prediction:
$\mathbf{x}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1},$ (9)
@@
which is more suitable for human motion generation. Detailed derivation is provided in the supplementary.
3.2 Physical Constraint Guidance
@@
There exist many physically implausible inter-penetrations between the actor and reactor in the generated results of current diffusion-based methods [69, 59, 15], owing to the stochastic sampling process that disregards physical constraints.
To address the issue, we propose a training-free physical constraint guidance method that calculates the penetration gradients $\nabla\mathcal{L}_{\text{pene}}$ that pulls the reactor in the opposite direction to the actor. Then these gradients are employed to guide the sampling process at each timestep to ensure the suppression of penetration. Specifically, We use $\psi^{h}_{i}(\mathbf{x})$ to represent the position of the $i$-th joint in the $h$-th frame of the generated reaction motion $\mathbf{x}$. Inspired by [30, 36], we calculate the signed distance function (SDF) between the actor and the reactor. Thus, our penetration loss functon is defined as:
@@
$\mathcal{L}_{\text{pene}}(\mathbf{x})\vcentcolon=\sum_{i,h}-\mathrm{min}\Bigl{(}\mathrm{SDF}(\psi^{h}_{i}(\mathbf{x})),\zeta\Bigr{)}$ (10)
@@
where $\zeta$ defines the safe distance between the actor and the reactor, beyond which the gradient becomes zero, and $\mathrm{SDF}$ is the signed distance function for an actor in the $h$-th frame, which dynamically changes across frames.
Flow maching sampling with physical constraint guidance. At each sampling step, the gradient of $\nabla\mathcal{L}_{\text{pene}}$ is utilized to guide the generative process. In contrast to the standard process for motion synthesis in Equation 8, the sampling process with physical constraint guidance is formalized as:
$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\mathbf{x}_{t_{n}}+(t_{n+1}-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (11)
@@
$\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1}),$ (12)
where $\mathbf{x}^{\prime}_{t_{n+1}}$ denoted the intermediate results of standard sampling in Equation 8. However, standard sampling based on $\mathbf{v}$-prediction often requires to estimate $\hat{\mathbf{x}}_{1}$:
$\hat{\mathbf{x}}_{1}\leftarrow\mathbf{x}_{t_{n}}+(1-t_{n})~{}\mathbf{v}_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c}),$ (13)
from current state $\mathbf{x}_{t_{n}}$ using an equivalent form of parameterization in Equation 7 to input into the signed distance function $\mathcal{L}_{\text{pene}}(\mathbf{x})$. In comparison, the sampling process based on $\mathbf{x}_{1}$-prediction with physical constraint guidance is as follows:
$\mathbf{x}^{\prime}_{t_{n+1}}\leftarrow\frac{1-t_{n+1}}{1-t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-t_{n}}~{}\hat{\mathbf{x}}_{1},$ (14)
$\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1}),$ (15)
@@
where $\hat{x}_{1}$ is the clean body poses predicted by a neural network $G_{\theta}$, which is more compatible with our physical constraint guidance method. The complete Euler sampling with physical constraint guidance for $\mathbf{x}_{1}$-prediction or $\mathbf{v}$-prediction is shown in Algorithm 1 and Algorithm 2 respectively. In practice, we use $\mathbf{x}_{1}$-prediction for its better performance. Under iterative sampling and physical constraint guidance, our method can generate more realistic and physically-plausible reaction motions.
@@
Our guidance method is actually a refined fine-tuning, which may be not suitable for training phases. In addition, the loss during the training phase mainly measures the difference between generated results and ground truth, while our guidance during the inference phase can provide more flexible guidance based on the quality of the generated results.
@@
Metrics. To qualitatively measure the degree of penetration, we introduced two metrics: 1) Intersection Volume (IV). Penetrate in [73, 21] just measures ground penetration which is not suitable for measuring the degree of penetration between humans. Interpenetration in [44] can only be computed as rigid bodies in the physics simulation. Inspired by Solid Intersection Volume (IV) [78, 41] in hand-object interaction, we measure human-human inter-penetration by voxelizing actor and reactor meshes and reporting the volume of voxels occupied by both. Intersection Volume (IV) is defined as
@@
where $V^{h}_{\text{pene}}$ represents the intersection volume of frame $h$ and $N_{\text{total}}$ denotes the total number of samples.
2) Intersection Frequency (IF). However, solely considering this metric can be misleading since it does not account for the case where there is no contact between the reactor and the reactor at all. Inspired by Contact Frequency in [36, 54], we introduce IF to measure the frequency of inter-penetration, i.e.
$IF=\frac{f_{\text{pene}}}{F_{\text{total}}}\cdot 100\%,$ (17)
where $f_{\text{pene}}$ represents the number of frames with inter-penetrations and $F_{\text{total}}$ denotes the total number of frames.
Algorithm 1 Euler Sampling algorithm with physical constraint guidance. ($\mathbf{x}_{1}$-prediction)
@@
6:     # Flow maching $\mathbf{x}_{1}$-prediction sampling (Eq. 14)
@@
8:     # Physical constraint guidance (Eq. 15)
9:     $\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}+\lambda_{\text{pene}}\nabla_{\mathbf{x}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
@@
11:  Return: The reaction motion after guidance $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$

### v3

@@ Abstract
Human action-reaction synthesis, a fundamental challenge in modeling causal human interactions, plays a critical role in applications ranging from virtual reality to social robotics. While diffusion-based models have demonstrated promising performance, they exhibit two key limitations for interaction synthesis: reliance on complex noise-to-reaction generators with intricate conditional mechanisms, and frequent physical violations in generated motions. To address these issues, we propose Action-Reaction Flow Matching (ARFlow), a novel framework that establishes direct action-to-reaction mappings, eliminating the need for complex conditional mechanisms. Our approach introduces a physical guidance mechanism specifically designed for Flow Matching (FM) that effectively prevents body penetration artifacts during sampling. Moreover, we discover the bias of traditional flow matching sampling algorithm and employ a reprojection method to revise the sampling direction of FM. To further enhance the reaction diversity, we incorporate randomness into the sampling process. Extensive experiments on NTU120, Chi3D and InterHuman datasets demonstrate that ARFlow not only outperforms existing methods in terms of Fréchet Inception Distance and motion diversity but also significantly reduces body collisions, as measured by our new Intersection Volume and Intersection Frequency metrics. Our project is available at https://arflow2025.github.io/.
@@ 1 Introduction
To address these challenges, we propose Action-Reaction Flow Matching (ARFlow), a novel framework that fundamentally resolves these limitations. Unlike diffusion models constrained by noise-data mappings, flow matching [40, 44, 1, 48] naturally models paired distributions through linear interpolation between endpoints (See Fig. 1), making it inherently suitable for action-reaction synthesis. ARFlow establishes direct pathways between action and reaction distributions, enabling simpler training and faster inference compared to diffusion-based methods. Instead of traditional vector field prediction ($\mathbf{v}$-prediction) in standard flow matching [40, 27], ARFlow directly outputs clean body poses ($\mathbf{x}_{1}$-prediction) in human motion generation. To eliminate unrealistic body collisions between characters, we further develop a specialized sampling algorithm that performs physical constraint guidance at $\mathbf{x}_{1}$ to avoid inaccurate gradients, and projects it back onto a revised Flow Matching path through interpolation. This innovation maintains physical plausibility through gradient guidance without compromising motion quality. Additionally, we introduce two new physics-aware evaluation metrics: Intersection Volume (IV) and Intersection Frequency (IF). Our main contributions are threefold:
@@
  • We integrate physical constraints and employ a reprojection method to revise the sampling direction of FM, surpassing vanilla guidance methods. This successfully prevents body collisions between characters while preserving natural movement. To further enhance the reaction diversity, we introduce randomness into the sampling process.
@@ 2.2 Human Action-Reaction Synthesis
Different from human-human interaction [70, 39, 58, 28, 60, 67], human action-reaction synthesis is causal and asymmetric [72, 42, 71]. To address this task, researchers have leveraged large language models [57, 59, 30], VAE-based methods [10, 45, 46]. However, these methods cannot capture fine-grained representations and ensure diversity, and diffusion-based methods [38, 60] are limited to the “offline” and “constrained” setting of human reaction generation, failing to generate instant and intention agnostic reactions. More recently, ReGenNet [72] introduce a diffusion-based transformer decoder framework and treat action sequence as conditional signal for online reaction generation. However, it often produce physically-implausible inter-penetrations between the actor and reactor since they disregard physical constraints in the generative process. Our method addresses this problem by ARFlow sampling with physical constraint guidance.
@@ 2.3 Flow Matching
Flow Matching [40, 44, 1, 48, 47, 17] has emerged as an efficient alternative to diffusion models, offering linear generation trajectories through ODE solvers. This paradigm enables simplified training and accelerated inference [16, 41], with successful applications spanning images [40, 16], audio [36], video [5], and point clouds [68]. In motion generation, MotionFlow [27] demonstrates comparable performance to diffusion models with faster sampling. Notably, Flow Matching inherently models transitions between arbitrary distributions through transport maps, making it particularly suitable for paired data modeling. Despite these advantages, its potential for action-reaction synthesis remains unexplored. Our work bridges this gap by establishing direct action-to-reaction mappings without complex conditional mechanisms.
@@ 3 Method
Figure 2: Pipeline of ARFlow. (a) For a sampled timestep $t$, we linearly interpolate a coupled action-reaction pair as Eq. 4 to produce the intermediate state $x^{1:H}_{t}$, which is then turns into a d-dimensional latent feature through a linear layer. We use Transformer Decoder Units to directly predict clean reaction motions. (b) After training the networks in (a), our ARFlow uses them for sampling, which is further guided by the gradient of $\mathcal{L}_{pene}$ to generate physically plausible reactions.
@@ 3.1 Human Action-Reaction Flow Matching
${~{}~{}~{}~{}\mathcal{L}_{\text{fm}}=\mathbb{E}_{\mathbf{x}_{1}\sim q(\mathbf{x}_{1}),\mathbf{x}_{0}\sim p(\mathbf{x}_{0}),t\sim[0,1]}[\|\mathbf{x}_{1}-G_{\theta}(\mathbf{x}_{t},t,c)\|_{2}^{2}].}$ (5)
@@
Since our neural network outputs $\hat{\mathbf{x}}_{1}$, we require to construct an equivalent relationship between the neural network’s predictions of $\mathbf{v}$ and $\mathbf{x}_{1}$. The equivalent form of parameterization Eq. 23 derived from our appendix is as follows:
${\mathbf{v}_{\theta}(\mathbf{x}_{t},t,c)=\frac{\hat{\mathbf{x}}_{1}-(1-\sigma_{\text{min}})\mathbf{x}_{t}}{1-(1-\sigma_{\text{min}})t},~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}}$ (7)
Then, our sampling based on $\mathbf{x}_{1}$-prediction can be achieved by first sampling $\mathbf{x}_{0}$ and then solving Eq. 1 employing an ODE solver [55, 35, 2] through our trained neural network $\hat{\mathbf{x}}_{1}=G_{\theta}(\mathbf{x}_{t},t,c)$. We use the Euler ODE solver and discretization process involves dividing the procedure into $N$ steps, leading to the following formulation:
@@
where the integer time step $t_{1}=0<t_{2}<\cdots<t_{N}=1$. By using equivalent form of parameterization Eq. 7, we finally obtain our flow maching sampling formulation based on $\mathbf{x}_{1}$-prediction:
$\mathbf{x}_{t_{n+1}}\leftarrow\frac{1-(1-\sigma_{\text{min}})t_{n+1}}{1-(1-\sigma_{\text{min}})t_{n}}\mathbf{x}_{t_{n}}+\frac{t_{n+1}-t_{n}}{1-(1-\sigma_{\text{min}})t_{n}}~{}\hat{\mathbf{x}}_{1},$ (9)
@@
which is more suitable for human motion generation. Detailed derivation is provided in Appendix A.
3.2 ARFlow Sampling with Physical Constraint Guidance
@@
There exist many physically implausible inter-penetrations between the actor and reactor in the generated results of current diffusion-based methods [72, 62, 15], owing to the sampling process that disregards physical constraints. To address the issue, we employ a penetration gradients $\nabla\mathcal{L}_{\text{pene}}$ to guide the sampling process. We follow [32, 38] to calculate the signed distance function (SDF) between the actor and the reactor as our penetration loss functon:
@@
$\mathcal{L}_{\text{pene}}(\mathbf{x})\vcentcolon=\sum_{i,h}-\mathrm{min}\Bigl{(}\mathrm{SDF}(\psi^{h}_{i}(\mathbf{x})),\zeta\Bigr{)},$ (10)
@@
where $\psi^{h}_{i}(\mathbf{x})$ represents the position of the $i$-th joint in the $h$-th frame of the generated reaction motion $\mathbf{x}$, the $\zeta$ defines the safe distance between the actor and the reactor, beyond which the gradient becomes zero, and $\mathrm{SDF}$ is the signed distance function for an actor in the $h$-th frame, which dynamically changes across frames.
Traditional diffusion guidance methods [12, 31, 63] first estimate $\hat{\mathbf{x}}_{1}$ from current state $\mathbf{x}_{t_{n}}$ with a denoiser network $\epsilon_{\theta}(\mathbf{x}_{t_{n}},t_{n},\mathbf{c})$, and then calculate gradients of the loss function with respect to current state $\mathbf{x}_{t_{n}}$, so it inevitably requires differentiation of the neural network, resulting in inaccurate gradients $\nabla\mathcal{L}_{\text{pene}}$. To avoid this issue, we directly updates the gradient at $\hat{\mathbf{x}}_{1}$:
@@
$\hat{\mathbf{x}}^{\prime}_{1}\leftarrow\hat{\mathbf{x}}_{1}-\lambda_{\text{pene}}\nabla_{\hat{\mathbf{x}}_{1}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1}),$ (11)
where $\hat{\mathbf{x}}_{1}$ is the clean body poses predicted by our neural network $G_{\theta}$ and $\lambda_{\text{pene}}$ is the guidance strength. Then, we use the linear interpolation of Flow Matching to project back to the intermediate state of learned FM path.
Interpolation between action and reaction. Since we build flow matching between the action and reaction distribution, unlike the process of adding noise back in diffusion, our interpolation process depends on the initial point $\mathbf{x}_{0}$, so we need to project towards the direction of the initial point. The Flow Matching sampling algorithm Eq. 9 is equivalent to the following formulation:
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\hat{\mathbf{x}}_{0}\leftarrow~{}$ $\hat{\mathbf{x}}_{1}+\frac{\mathbf{x}_{t_{n}}-(1+\sigma_{\text{min}}t)\hat{\mathbf{x}}_{1}}{1-(1-\sigma_{\text{min}})t_{n}},$ (12)
$\mathbf{x}_{t_{n+1}}\leftarrow~{}$ $t_{n+1}\hat{\mathbf{x}}_{1}+[1-(1-\sigma_{\text{min}})t_{n+1}]~{}\hat{\mathbf{x}}_{0}.$ (13)
This sampling process essentially finds a $\hat{\mathbf{x}}_{0}$ along the opposite direction of the current velocity field for linear interpolation as Eq. 12 (black dotted lines in Fig. 3). Obviously, this predicted $\hat{\mathbf{x}}_{0}$ deviates from the initial point $\mathbf{x}_{0}$ (purple dotted lines in Fig. 3). In order to correct the projection direction of traditional Flow Matching, we use a weight factor $w$ to weight $\hat{\mathbf{x}}_{0}$ and $\mathbf{x}_{0}$:
$\hat{\mathbf{x}}^{*}_{0}\leftarrow w\hat{\mathbf{x}}_{0}+(1-w)\mathbf{x}_{0}.$ (14)
@@
and use $\hat{\mathbf{x}}^{*}_{0}$ as our final endpoint for interpolation. The ARFlow sampling and traditional sampling algorithm for $\mathbf{x}_{1}$-prediction with physical constraint guidance are shown in Algorithm 2 and 1 respectively. In practice, we use $\mathbf{x}_{1}$-prediction for its better performance. Under iterative sampling and physical constraint guidance, our method can generate more realistic and physically-plausible reaction motions.
Stochastic sampling to enhance diversity of reactions. To further enhance the diversity of reactions generated by ARFlow, we incorporate randomness into the sampling process. The interpolation Eq. 13 can be written in the following equivalent form:
$\mathbf{x}_{t_{n+1}}\leftarrow\hat{\mathbf{x}}_{1}+(1-t_{n+1})(\hat{\mathbf{x}}_{0}-\hat{\mathbf{x}}_{1})+\sigma_{\text{min}}t_{n+1}\hat{\mathbf{x}}_{0}.$ (15)
The interpolation process can be understood as a projection in the opposite direction of the current learned velocity field $\hat{\mathbf{x}}_{1}-{\hat{\mathbf{x}}}_{0}$. Thus, we can weight the projection direction $\hat{\mathbf{x}}^{*}_{0}-\hat{\mathbf{x}}^{\prime}_{1}$ and stochastic direction $d_{\text{random}}$ to incorporate randomness:
$d_{\text{mix}}\leftarrow\hat{\mathbf{x}}^{*}_{0}-\hat{\mathbf{x}}^{\prime}_{1}+\beta[d_{\text{random}}-(\hat{\mathbf{x}}^{*}_{0}-\hat{\mathbf{x}}^{\prime}_{1})],$ (16)
$\mathbf{x}_{t_{n+1}}\leftarrow~{}\hat{\mathbf{x}}^{\prime}_{1}+(1-t_{n+1})d_{\text{mix}}+\sigma_{\text{min}}t_{n+1}\hat{\mathbf{x}}^{*}_{0},$ (17)
@@
where $\beta$ is the factor to control the strength of randomness. Our guidance method is actually a refined fine-tuning, which may be not suitable for training. In addition, the loss during the training mainly measures the difference between generated results and ground truth, while our guidance during the inference phase can provide more flexible guidance based on the quality of the generated results.
Figure 3: Illustration of our ARFlow sampling with physical constraint guidance.
FID $\downarrow$ IF $\downarrow$ IV $\downarrow$
Real 0.09 21.96% 5.35
       Offline
MDM 22.04 15.45% 3.36
ReGenNet 6.19 19.14% 3.07
ARFlow 5.00 9.67% 1.94
MDM ${}^{\text{vanilla}}$ 20.42 13.29% 2.26
ReGenNet ${}^{\text{vanilla}}$ 7.69 17.85% 2.16
[…]
ARFlow ${}^{\text{improved}}$ 5.05 3.73% 0.23
       Online
MDM 51.11 32.63% 17.97
ReGenNet 11.00 13.84% 3.50
ARFlow 7.89 8.39% 3.26
MDM ${}^{\text{vanilla}}$ 40.03 22.63% 10.35
ReGenNet ${}^{\text{vanilla}}$ 11.63 13.05% 2.71
ARFlow ${}^{\text{vanilla}}$ 8.61 3.29% 1.44
ARFlow ${}^{\text{improved}}$ 8.07 3.23% 0.53
Metrics. To qualitatively measure the degree of penetration, we introduced two metrics:
@@
1) Intersection Volume (IV). Penetrate in [76, 23] just measures ground penetration which is not suitable for measuring the degree of penetration between humans. Interpenetration in [46] can only be computed as rigid bodies in the physics simulation. Inspired by Solid Intersection Volume (IV) [81, 43], we measure human-human inter-penetration by voxelizing actor and reactor meshes and reporting the volume of voxels occupied by both. Intersection Volume (IV) is defined as
@@
where $V^{h}_{\text{pene}}$ represents intersection volume of frame $h$ and $N_{\text{total}}$ denotes the total number of samples.
2) Intersection Frequency (IF). Inspired by Contact Frequency in [38, 57], we introduce IF to measure the frequency of inter-penetration, i.e.
$IF=f_{\text{pene}}/F_{\text{total}},$ (19)
where $f_{\text{pene}}$ represents the number of inter-penetration frames and $F_{\text{total}}$ is the total number of frames.
Algorithm 1 Sampling algorithm with vanilla guidance of physical constraints.
@@
2:  Parameters: $N$ the number of sampling steps; $\lambda_{\text{pene}}$ the guidance strength
@@
6:     # Flow Matching $\mathbf{x}_{1}$-prediction sampling (Eq. 9)
@@
8:     # Physical constraint guidance
9:     $\mathbf{x}_{t_{n+1}}\leftarrow\mathbf{x}^{\prime}_{t_{n+1}}-\lambda_{\text{pene}}\nabla_{\hat{\mathbf{x}}_{t_{n}}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
@@
11:  Return: The reaction motion $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$
Algorithm 2 Sampling algorithm with improved guidance of physical constraints.
2:  Parameters: $N$ the number of sampling steps; $\lambda_{\text{pene}}$ the guidance strength; $w$ weight factor
3:  Sample $\mathbf{x}_{0}$ from the action distribution
4:  for $n=1,2,...,N-1$ do
6:     $\hat{\mathbf{x}}_{0}\leftarrow\hat{\mathbf{x}}_{1}+\frac{\mathbf{x}_{t_{n}}-(1+\sigma_{\text{min}}t)\hat{\mathbf{x}}_{1}}{1-(1-\sigma_{\text{min}})t_{n}}$       #  (Eq. 12)
7:     # Physical constraint guidance at $\hat{\mathbf{x}}_{1}$(Eq. 11)
8:     $\hat{\mathbf{x}}^{\prime}_{1}\leftarrow\hat{\mathbf{x}}_{1}-\lambda_{\text{pene}}\nabla_{\hat{\mathbf{x}}_{1}}\mathcal{L}_{\text{pene}}(\hat{\mathbf{x}}_{1})$
9:     # Direction correction
10:     $\hat{\mathbf{x}}^{*}_{0}\leftarrow w\hat{\mathbf{x}}_{0}+(1-w)\mathbf{x}_{0}$
11:     # Interpolation (Eq. 13)
12:     $\mathbf{x}_{t_{n+1}}\leftarrow t_{n+1}\hat{\mathbf{x}}^{\prime}_{1}+[1-(1-\sigma_{\text{min}})t_{n+1}]~{}\hat{\mathbf{x}}^{*}_{0}$
13:  end for
14:  Return: The reaction motion $\mathbf{x}_{1}=\mathbf{x}_{t_{N}}$